OpenStudy (rosemarybaldwin7):

The table below shows some input and output values for a function named F(n). Which of the following equations describes F(n)? N F(n) 0 5 1 8 2 11 5 20 10 35 20 65 a:F(n)=4n b:F(n)=n+5 c:F(n)=3n+5

OpenStudy (texaschic101):

just a matter of subbing... f(n) = 3n + 5 now sub.. 5 = 3(0) + 5......5 = 5...correct and u will find that when using f(n) = 3n + 5 and subbing in ur inputs and outputs...they will all equal.